OpenStudy (anonymous):

a mummy is found to be 3940 years old by ^14C dating. Using half life of 5730 years for ^14C demonstrate that 62.1% of ^14C at the death of the person remains the mummy

OpenStudy (aaronq):

For 1st order reactions you can use these formulas: find the decay constant \[t _{1/2}=\frac{ \ln2 }{ k }\] t1/2 = half-life then use the general exponential decay/growth equation: \[A _{t}=A _{0}*e ^{-kt}\] Ao=initial amount At= amount after time elapsed t=time elapsed k=decay constant start off by assuming you have 100% 14C, i.e. Ao=100

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

.125=e^-k(5730)

OpenStudy (aaronq):

you set it up wrong

OpenStudy (aaronq):

find k first

OpenStudy (anonymous):

how?

OpenStudy (aaronq):

with the first equation i wrote

OpenStudy (anonymous):

what would At be though

OpenStudy (aaronq):

okay rewind a little bit, its asking you to prove that 62.1% remains of the mummy. now you can do this several ways, find what At is assuming all other variables OR use 62.1% for At and find the initial amount Ao.

OpenStudy (anonymous):

im sorry I still don't understand if Ao equals 100 then what would At and t be

OpenStudy (aaronq):

well if you're using Ao as 100% then you're solving for At if you're using At as 62.1% then you're solving for Ao

OpenStudy (aaronq):

t is time elapsed so 3940 years

OpenStudy (anonymous):

what is the k value we are using

OpenStudy (aaronq):

you found it in the first equation using the half life

OpenStudy (anonymous):

ok then I got 1.21x10^-4 for k

OpenStudy (aaronq):

use it in the other equation

OpenStudy (anonymous):

i did and i got .621 which is 62.1% right?

OpenStudy (aaronq):

yep, in decimal, 0.621 is 62.1%

OpenStudy (anonymous):

ok thanks so much do you have time for one more?

OpenStudy (aaronq):

no problem, yeah, what is it?

OpenStudy (anonymous):

nevermind i figured it out thanks so much i was so screwed because i have test tomorrow

OpenStudy (aaronq):

okay, good luck on your test !